# arXiv:2603.03190v1[cs.AI] 03 Mar 2026

### v1

@@ 3 Discussion
We next discuss limitations and future directions, focusing on Surprisal/Entropy computation and dataset requirements. In this study, we used MusicGen [musicgen] as the music language model to compute Surprisal and Entropy directly from audio. Our first criterion in selecting the model was that it should reflect long-term statistical regularities acquired from a large-scale music corpus, because expectation features derived from such long-term learning better explain cortical responses than those based only on short-term regularities [Kern2023]. In this line of work, AudioLM established the general framework of discrete audio language modeling without symbolic representations, while Jukebox and MusicLM provided music-specific raw-audio alternatives trained at large scale [AudioLM, Jukebox2020, MusicLM2023]. We selected MusicGen because it offers a single-stage autoregressive Transformer over EnCodec tokens, rather than a hierarchical multi-stage generation pipeline [musicgen, encodec]. This makes next-token probabilities straightforward to extract from one model over one tokenization scheme, while still leveraging a large licensed music corpus and strong music-generation quality [musicgen]. For these reasons, MusicGen provided the most practical and conceptually clean basis for computing frame-wise Surprisal and Entropy in the present study.
@@
MusicGen builds an autoregressive model over four discrete EnCodec Residual Vector Quantization (RVQ) codebooks per time step [encodec]. Let $q_{t}^{(k)}$ denote the token at physical time $t$ in codebook $k\in\{1,2,3,4\}$. In principle, the Surprisal of the full EnCodec representation at physical time $t$ is the joint quantity
@@
where $Q_{<t}$ denotes all RVQ tokens at earlier physical times. We do not compute this joint quantity because the publicly released MusicGen model is trained with the delay interleaving pattern, which is computationally efficient but yields an inexact factorization across codebooks [musicgen]. This choice leads to two practical considerations. First, under the delay pattern, multiple codebooks are predicted in parallel at each sequence step. Consequently, when the model assigns a probability to the $k1$ token $q_{t}^{(1)}$, its conditioning context $C_{t}^{\mathrm{delay}}$ can exclude some residual-stream tokens from the recent past (e.g., $q_{t-1}^{(2)}$), because those tokens are predicted in the same step. We therefore compute Surprisal and Entropy only for codebook 1 ($k1$) using the native conditional distribution $p_{\theta}\!\left(q_{t}^{(1)}\mid C_{t}^{\mathrm{delay}}\right)$. When interpreting these values, we treat them as a proxy for the conditional that would be obtained under an exact flattening factorization with the full past context,
@@
where $C_{t}^{\mathrm{full}}$ denotes all past RVQ tokens across codebooks. We judge this approximation operationally acceptable because delay-pattern MusicGen achieves high-quality generation [musicgen]. Second, when predicting codebooks $k3$ and $k4$, the delay pattern can introduce future references, as the conditioning context may contain information from later time steps in other codebooks. Because our goal is to approximate brain-like predictive probabilities, such future reference is undesirable. Therefore, predictive quantities that avoid future information leakage are limited to $k1$ (and potentially $k2$). However, since $k2$ at time $t$ depends on $k1$ at the same physical time step, using $k2$ would violate strict causal interpretation. To preserve causal validity of expectation features, we therefore restrict computation to $k1$ only. In RVQ, each quantizer encodes the residual error of the previous stage, and prior RVQ-based work characterizes the first codebook as the most important component of the representation [musicgen, encodec]. Focusing on $k1$ therefore targets the primary signal structure.
As a complementary direction, future work could employ an acoustic language model trained with an exact flattening factorization across all RVQ codebooks. This would allow computation of joint probabilities over the full multi-codebook representation without relying on the delay-pattern approximation. Furthermore, if future studies establish that different RVQ codebooks capture partially disentangled musical attributes such as melody, rhythm, harmony, dynamics, or timbre [Wang2022DRL], then codebook-specific Surprisal and Entropy could be interpreted as attribute-specific predictive quantities, enabling finer-grained analysis of musical expectation.
@@ Methods
The primary objective of the pretraining stage is to guide the encoder to extract EEG representations that preserve stimulus-related information encoded in music-derived teacher signals. Specifically, given a 3-s EEG segment $x$, the encoder–decoder pathway is trained such that the decoded representation $G(F(x))$ can predict masked tokens of a discretized music feature sequence $m$ computed from the corresponding audio stimulus. Inspired by the Supervised Masked Autoencoders (SupMAE) framework [supmae], Song ID classification from $F(x)$ is incorporated concurrently as an auxiliary objective. While the nominal loss weight for the classification task is set higher than that of the masked prediction task, it is important to note that the prediction loss is computed over a dense sequence of numerous masked tokens per segment. Consequently, frame-level predictive signal dominates the overall learning dynamics. Sequence-level classification objective serves to guide and stabilize the representation learning process rather than acting as the primary driver of pretraining.
The teacher signal $m$ always consists of discretized music features. These features are obtained from two distinct audio-based models: (i) MuQ features derived from a masked language modeling-based self-supervised audio model and discretized using $k$-means clustering, and (ii) Surprisal and Entropy values computed from an autoregressive audio language model and discretized using quantile binning (see Fig. 9, Supplementary Note 1 and Supplementary Note 2 for details of the processing pipeline).
@@
Our framework is designed to integrate these diverse representations, leveraging the finding that the human brain encodes both acoustic [Bellier2023, Tuckute2023, Mischler2025] and expectation [DiLiberto2020, Kern2023, GaleanoOtalaro2024] information. Since the inclusion of expectation features has been shown to improve the prediction of neural responses beyond what is achieved by acoustic features alone [DiLiberto2020], we employed an ensemble strategy to combine encoders pretrained with these different teacher signals. By integrating these complementary representations, our approach aims to capture the diverse information axes, thereby enhancing the performance and robustness of EEG-based song ID classification (see Discussion for a detailed analysis of ensemble effects).
@@
We used the Naturalistic Music EEG Dataset–Tempo (NMED-T) [losorelli2017nmed], a publicly available dataset consisting of EEG recordings acquired while participants listened to naturalistic music. The dataset includes EEG data from 20 participants who listened to 10 commercially released musical pieces. Neural activity was recorded from 128 scalp electrodes during continuous music listening.
EEG signals in the original dataset were recorded at a sampling rate of 1000 Hz. Following the preprocessing pipeline provided by Losorelli et al. [losorelli2017nmed], the raw EEG signals were high-pass filtered at 0.3 Hz, notch filtered at 59–61 Hz to remove line noise, and low-pass filtered at 50 Hz. After filtering, the signals were downsampled to 125 Hz. We used the preprocessed EEG signals provided by the dataset without modifying the original filter settings, ensuring consistency with prior analyses of the NMED-T dataset.
Unless otherwise noted, the dataset configuration, data splits, and EEG preprocessing procedures used in this study closely follow those described in Akama et al. [Akama2025]. This design choice facilitates direct and controlled comparisons of Song ID classification performance between the proposed framework and the original PredANN method described therein.
@@
All EEG signals were analyzed at a sampling rate of 125 Hz with 128 scalp channels, following the preprocessing provided with the NMED-T dataset [losorelli2017nmed]. To ensure consistency across recordings, all EEG recordings were truncated to a common maximum duration of 4 min. Each 4-min recording was then segmented into non-overlapping 30-s excerpts, resulting in eight excerpts per musical piece.
The 30-s excerpts were divided into training and validation sets using a 75:25 split. To preserve the original song distribution in both sets, stratified sampling was applied with respect to Song ID labels. Specifically, we used the train_test_split function from the sklearn.model_selection module, specifying the stratify parameter as the song label [scikit]. To guarantee reproducibility, the random seed was fixed to 42, consistent with the experimental setup of Akama et al. [Akama2025].
Based on prior neurophysiological findings regarding temporal delays in auditory cortical responses reflecting melodic expectations [DiLiberto2020], a fixed temporal shift of 200 ms was applied to align EEG signals with the corresponding auditory stimuli. Concretely, the starting point of each EEG segment was shifted forward by 200 ms, which corresponds to 25 samples at a sampling rate of 125 Hz. This temporal alignment is consistent with evidence that higher-order auditory cortical responses to music typically emerge with latencies in the range of approximately 150–250 ms [DiLiberto2020]. Importantly, this delay value was not tuned in the present study, but was adopted directly from Akama et al. [Akama2025], who systematically compared multiple latency conditions and demonstrated that a delay of approximately 200 ms yielded the highest song ID classification accuracy.
@@
EEG normalization was performed independently for each 3-s EEG segment and for each channel. Specifically, a RobustScaler [scikit] based on the median and interquartile range was applied to the time series of each channel, followed by clamping the scaled values to the range [20, 20][-20,\,20]. This normalization strategy suppresses the influence of transient large-amplitude noise and artifacts while preserving the relative temporal structure of the EEG signals. Because normalization was applied independently to each sample, no statistics were shared across training, validation, or test sets, thereby preventing information leakage. Implementation details are provided in the function normalize_EEG_4 in the file preprocessing_eegmusic_dataset_3s.py.
@@
We constructed three types of artificial neural network (ANN) representations directly from the acoustic signals, as summarized in Figure 9. These representations were designed to capture complementary aspects of auditory processing and were used as teacher signals for masked-prediction-based multitask pretraining. Specifically, we extracted (i) an acoustic representation derived from a self-supervised music foundation model [zhu2025muqselfsupervisedmusicrepresentation], and (ii) predictive representations derived from Surprisal and Entropy computed from an autoregressive music language model [musicgen]. All three ANN representations were discretized into 128 levels (0–127) and served as prediction targets during pretraining.
In the PredANN framework, a central assumption is that ANN representations that closely resemble human cortical representations can serve as effective teacher signals for EEG-based representation learning [Akama2025]. Accordingly, we aimed to use ANN representations that are as neurophysiologically plausible as possible. A growing body of prior work has demonstrated that self-supervised learning (SSL) models capture human cortical representations more faithfully than supervised models, particularly in the auditory domain [NEURIPS2022_d81ecfc8, vaidya2022selfsupervisedmodelsaudioeffectively, oota2023, oota2024deepneuralnetworksbrain]. These findings motivate the use of SSL-based models for extracting teacher representations from audio.
Moreover, prior studies have shown that the architectural properties of ANN models play a critical role in their correspondence with cortical representations. In particular, convolutional layers have been reported to exhibit lower predictive power for brain activity compared with Transformer layers [NEURIPS2022_d81ecfc8]. In contrast, the functional hierarchy formed by Transformer layers has been shown to align well with the hierarchical organization of the human cerebral cortex [NEURIPS2022_d81ecfc8]. Consistent with these observations, multiple studies suggest that Transformer-based models exhibit stronger representational similarity to cortical responses than alternative architectures [NEURIPS2022_d81ecfc8, vaidya2022selfsupervisedmodelsaudioeffectively].
Based on this converging evidence, we adopted Transformer-based SSL models to extract ANN representations from audio. Specifically, we used MuQ [zhu2025muqselfsupervisedmusicrepresentation], a masked-prediction-based Transformer model, to obtain acoustic representations, and MusicGen [musicgen], an autoregressive Transformer model, to derive predictive representations in the form of Surprisal and Entropy. The detailed procedures for computing and discretizing these representations are described in the following sections and formalized in dedicated algorithms in Supplementary Note 1 and Supplementary Note 2.
@@
To construct ANN representations that primarily encode acoustic information, we employed MuQ [zhu2025muqselfsupervisedmusicrepresentation], a Transformer-based masked-prediction self-supervised music foundation model. MuQ has been shown to outperform representative SSL models such as MERT [mert] and MusicFM [musicfm] across nearly all tasks in the MARBLE benchmark [marble]. Additionally, MuQ-MuLan, which incorporates MuQ as an audio encoder, achieves state-of-the-art performance in zero-shot music tagging, surpassing MuLan [mulan2022] and Microsoft-CLAP 2023 [elizalde2024clap]. These results indicate that MuQ provides a strong and expressive acoustic representation suitable as a teacher signal.
For each musical piece, continuous MuQ embeddings were extracted from 30-second audio chunks following the original MuQ preprocessing configuration. Audio signals were converted to mono by channel averaging, segmented into 30-second waveforms, and resampled to 24 kHz using a Kaiser-windowed sinc filter, implemented in the librosa package [mcfee_librosa] via librosa.resample with res_type="kaiser_fast". MuQ was then applied to each 30-second chunk, and the last-layer hidden representations were extracted, yielding a sequence of frame-wise embeddings at a temporal resolution of 25 Hz (one embedding every 40 ms). Formally, let $\mathcal{T}_{30\mathrm{s}}=\{1,\dots,1500\}$ denote the set of time frame indices within a 30-second audio chunk, where the temporal resolution is 25 Hz (40-ms intervals). The resulting continuous acoustic representation $\mathbf{M}_{\mathrm{raw}}$ is defined as:
@@
The use of 24 kHz sampling rate and fixed 30-second input length strictly follows the MuQ pretraining setup [zhu2025muqselfsupervisedmusicrepresentation]. Extracting features under conditions that closely match the pretraining distribution reduces out-of-distribution effects and ensures that the resulting embeddings faithfully reflect the learned acoustic structure.
@@
Thus, the model explicitly distinguishes between continuous acoustic features $\mathbf{m}_{\mathrm{raw}_{t}}$ used as decoder inputs and discrete acoustic tokens $m_{\mathrm{disc}_{t}}$ used as supervision signals, while maintaining a one-to-one temporal correspondence between them. This frame-wise formulation provides the necessary foundation for constructing 3-second MuQ representations aligned with EEG segments in subsequent stages of the model. The detailed procedures for extracting continuous MuQ embeddings and performing K-means discretization are formalized in Supplementary Note 1 (Algorithm 1 and Algorithm 2).
@@
To construct ANN representations that primarily encode predictive information, we employed MusicGen-large (facebook/musicgen-large) from the Audiocraft framework [musicgen] as an autoregressive model. MusicGen is the first model to achieve high-quality, long-duration, and controllable music generation using a single-stage Transformer language model. It is trained as a language model over discrete audio tokens produced by EnCodec [encodec], which applies residual vector quantization (RVQ) with multiple codebooks.
@@
A key contribution of MusicGen lies in its treatment of multi-codebook RVQ tokens through codebook interleaving patterns. Copet et al. proposed and evaluated several patterns, including flattening, delay, parallel, and coarse first [musicgen]. While the flattening pattern provides exact autoregressive factorization, the delay pattern uses an inexact but computationally efficient decomposition. Despite reducing computational cost by approximately a factor of four, evaluation showed that the delay pattern achieves similar generation quality to flattening. The MusicGen-large model used in this study is trained with the delay pattern, which we adopt throughout our experiments. We compute predictive features under an empty text condition, corresponding to the model’s unconditional distribution learned through classifier-free guidance training.
@@
The continuous frame-wise Surprisal and Entropy sequences, referred to as Raw Surprisal and Raw Entropy, are used as continuous decoder inputs during masked-prediction pretraining. In parallel, these continuous values are discretized into 128 levels using equal-frequency (quantile) binning to obtain discrete predictive tokens used as ground-truth supervision. Specifically, for each feature type, all values across all songs and segments are pooled to form a one-dimensional set $u$. The global quantile boundaries are then computed as:
$q_{k}=\mathrm{Quantile}(u,k/128),\quad k=0,\dots,127$
where $q_{k}$ represents the threshold below which a fraction $k/128$ of the values in $u$ falls. Each value is then assigned a bin index in $\{0,\dots,127\}$ using these global boundaries. By replacing each continuous value with its corresponding discrete token, the resulting discrete sequences (Discretized Surprisal and Discretized Entropy) maintain a strict one-to-one temporal correspondence with the continuous inputs.
The complete procedures for computing frame-wise Surprisal and Entropy and for quantile-based discretization are formalized in Supplementary Note 2 (Algorithm 1 and Algorithm 2).
@@
Specifically, for each 30-second chunk, MusicGen [musicgen] is applied to compute k1 logits at a temporal resolution of 20 ms, yielding 1500 frames per chunk. Surprisal and Entropy are then computed frame-wise from these logits, and subsequently discretized into 128 bins using the same quantile-based discretization procedure as in the default setting. By construction, this approach prevents MusicGen from exploiting autoregressive context that spans across adjacent 30-second chunks, thereby strictly eliminating potential information leakage from unused acoustic context. The results of this conservative computation are presented in Supplementary Fig. 1.
@@
For MuQ [zhu2025muqselfsupervisedmusicrepresentation], which provides acoustic embeddings at a temporal resolution of 40 ms (25 Hz), the starting frame index is defined as
@@
Recent progress in self-supervised pretraining has established masked modeling with Transformer architectures as a powerful paradigm for representation learning. In natural language processing, BERT [devlin-etal-2019-bert] demonstrated that predicting masked tokens enables the acquisition of rich bidirectional contextual representations, while the GPT series showed that autoregressive pretraining effectively captures long-range dependencies [GPT, GPT2]. In computer vision, this idea was reformulated as masked image modeling, where approaches such as MAE [mae], BEiT [beit], and SimMIM [xie2021simmim] revealed that reconstructing missing content from partial observations leads to highly transferable representations.
A key step in transferring these ideas to EEG representation learning is exemplified by LaBraM [jiang2024large]. Inspired by masked modeling in vision, it reformulates EEG signals as collections of time–channel patches and applies Transformer-based masked reconstruction to large-scale EEG data. Through architectural choices such as temporal patch encoders, learnable channel and time embeddings, and Pre-Norm self-attention, LaBraM demonstrated that masked modeling is an effective and scalable principle for EEG representation learning.
The original PredANN framework, by contrast, was built around a CLIP [clip]-style contrastive learning objective using CNN-based encoders optimized with an InfoNCE loss. While this design successfully aligned EEG representations with stimulus-derived ANN representations, it fundamentally differs from the generative and structural learning paradigm underlying recent masked autoencoding approaches. Motivated by the success of LaBraM and masked modeling in other modalities, we therefore redesign PredANN into a Transformer-based masked modeling framework.
Specifically, we propose PredANN++, which extends PredANN by replacing the CLIP-style CNN encoder with a masked autoencoding Transformer architecture. Rather than discretizing EEG signals themselves as in LaBraM, PredANN++ treats discretized stimulus-derived representations (MuQ [zhu2025muqselfsupervisedmusicrepresentation], Surprisal, and Entropy) as teacher sequences and formulates pretraining as masked prediction of these discrete targets. This design preserves the original PredANN philosophy of learning EEG representations aligned with stimulus-level ANN representations, while adopting the masked generative learning paradigm shown to be effective in LaBraM. The overall architecture is illustrated in Fig. 8 and is formalized below.

### v2

@@ 3 Discussion
We next discuss limitations and future directions, focusing on Surprisal/Entropy computation and dataset requirements. In this study, we used MusicGen [43] as the music language model to compute Surprisal and Entropy directly from audio. Our first criterion in selecting the model was that it should reflect long-term statistical regularities acquired from a large-scale music corpus, because expectation features derived from such long-term learning better explain cortical responses than those based only on short-term regularities [32]. In this line of work, AudioLM established the general framework of discrete audio language modeling without symbolic representations, while Jukebox and MusicLM provided music-specific raw-audio alternatives trained at large scale [51, 52, 53]. We selected MusicGen because it offers a single-stage autoregressive Transformer over EnCodec tokens, rather than a hierarchical multi-stage generation pipeline [43, 46]. This makes next-token probabilities straightforward to extract from one model over one tokenization scheme, while still leveraging a large licensed music corpus and strong music-generation quality [43]. For these reasons, MusicGen provided the most practical and conceptually clean basis for computing frame-wise Surprisal and Entropy in the present study.
@@
MusicGen builds an autoregressive model over four discrete EnCodec Residual Vector Quantization (RVQ) codebooks per time step [46]. Let $q_{t}^{(k)}$ denote the token at physical time $t$ in codebook $k\in\{1,2,3,4\}$. In principle, the Surprisal of the full EnCodec representation at physical time $t$ is the joint quantity
@@
where $Q_{<t}$ denotes all RVQ tokens at earlier physical times. We do not compute this joint quantity because the publicly released MusicGen model is trained with the delay interleaving pattern, which is computationally efficient but yields an inexact factorization across codebooks [43]. This choice leads to two practical considerations. First, under the delay pattern, multiple codebooks are predicted in parallel at each sequence step. Consequently, when the model assigns a probability to the $k1$ token $q_{t}^{(1)}$, its conditioning context $C_{t}^{\mathrm{delay}}$ can exclude some residual-stream tokens from the recent past (e.g., $q_{t-1}^{(2)}$), because those tokens are predicted in the same step. We therefore compute Surprisal and Entropy only for codebook 1 ($k1$) using the native conditional distribution $p_{\theta}\!\left(q_{t}^{(1)}\mid C_{t}^{\mathrm{delay}}\right)$. When interpreting these values, we treat them as a proxy for the conditional that would be obtained under an exact flattening factorization with the full past context,
@@
where $C_{t}^{\mathrm{full}}$ denotes all past RVQ tokens across codebooks. We judge this approximation operationally acceptable because delay-pattern MusicGen achieves high-quality generation [43]. Second, when predicting codebooks $k3$ and $k4$, the delay pattern can introduce future references, as the conditioning context may contain information from later time steps in other codebooks. Because our goal is to approximate brain-like predictive probabilities, such future reference is undesirable. Therefore, predictive quantities that avoid future information leakage are limited to $k1$ (and potentially $k2$). However, since $k2$ at time $t$ depends on $k1$ at the same physical time step, using $k2$ would violate strict causal interpretation. To preserve causal validity of expectation features, we therefore restrict computation to $k1$ only. In RVQ, each quantizer encodes the residual error of the previous stage, and prior RVQ-based work characterizes the first codebook as the most important component of the representation [43, 46]. Focusing on $k1$ therefore targets the primary signal structure.
As a complementary direction, future work could employ an acoustic language model trained with an exact flattening factorization across all RVQ codebooks. This would allow computation of joint probabilities over the full multi-codebook representation without relying on the delay-pattern approximation. Furthermore, if future studies establish that different RVQ codebooks capture partially disentangled musical attributes such as melody, rhythm, harmony, dynamics, or timbre [54], then codebook-specific Surprisal and Entropy could be interpreted as attribute-specific predictive quantities, enabling finer-grained analysis of musical expectation.
@@ Methods
The primary objective of the pretraining stage is to guide the encoder to extract EEG representations that preserve stimulus-related information encoded in music-derived teacher signals. Specifically, given a 3-s EEG segment $x$, the encoder–decoder pathway is trained such that the decoded representation $G(F(x))$ can predict masked tokens of a discretized music feature sequence $m$ computed from the corresponding audio stimulus. Inspired by the Supervised Masked Autoencoders (SupMAE) framework [55], Song ID classification from $F(x)$ is incorporated concurrently as an auxiliary objective. While the nominal loss weight for the classification task is set higher than that of the masked prediction task, it is important to note that the prediction loss is computed over a dense sequence of numerous masked tokens per segment. Consequently, frame-level predictive signal dominates the overall learning dynamics. Sequence-level classification objective serves to guide and stabilize the representation learning process rather than acting as the primary driver of pretraining.
The teacher signal $m$ always consists of discretized music features. These features are obtained from two distinct audio-based models: (i) MuQ features derived from a masked language modeling-based self-supervised audio model and discretized using $k$-means clustering, and (ii) Surprisal and Entropy values computed from an autoregressive audio language model and discretized using quantile binning (see Fig. 9, Supplementary Note 1: Algorithmic details of MuQ extraction and Supplementary Note 2: Algorithmic details of computing Surprisal, Entropy, and discretization for details of the processing pipeline).
@@
Our framework is designed to integrate these diverse representations, leveraging the finding that the human brain encodes both acoustic [35, 36, 34] and expectation [14, 32, 33] information. Since the inclusion of expectation features has been shown to improve the prediction of neural responses beyond what is achieved by acoustic features alone [14], we employed an ensemble strategy to combine encoders pretrained with these different teacher signals. By integrating these complementary representations, our approach aims to capture the diverse information axes, thereby enhancing the performance and robustness of EEG-based song ID classification (see Discussion for a detailed analysis of ensemble effects).
@@
We used the Naturalistic Music EEG Dataset–Tempo (NMED-T) [41], a publicly available dataset consisting of EEG recordings acquired while participants listened to naturalistic music. The dataset includes EEG data from 20 participants who listened to 10 commercially released musical pieces. Neural activity was recorded from 128 scalp electrodes during continuous music listening.
EEG signals in the original dataset were recorded at a sampling rate of 1000 Hz. Following the preprocessing pipeline provided by Losorelli et al. [41], the raw EEG signals were high-pass filtered at 0.3 Hz, notch filtered at 59–61 Hz to remove line noise, and low-pass filtered at 50 Hz. After filtering, the signals were downsampled to 125 Hz. We used the preprocessed EEG signals provided by the dataset without modifying the original filter settings, ensuring consistency with prior analyses of the NMED-T dataset.
Unless otherwise noted, the dataset configuration, data splits, and EEG preprocessing procedures used in this study closely follow those described in Akama et al. [37]. This design choice facilitates direct and controlled comparisons of Song ID classification performance between the proposed framework and the original PredANN method described therein.
@@
All EEG signals were analyzed at a sampling rate of 125 Hz with 128 scalp channels, following the preprocessing provided with the NMED-T dataset [41]. To ensure consistency across recordings, all EEG recordings were truncated to a common maximum duration of 4 min. Each 4-min recording was then segmented into non-overlapping 30-s excerpts, resulting in eight excerpts per musical piece.
The 30-s excerpts were divided into training and validation sets using a 75:25 split. To preserve the original song distribution in both sets, stratified sampling was applied with respect to Song ID labels. Specifically, we used the train_test_split function from the sklearn.model_selection module, specifying the stratify parameter as the song label [56]. To guarantee reproducibility, the random seed was fixed to 42, consistent with the experimental setup of Akama et al. [37].
Based on prior neurophysiological findings regarding temporal delays in auditory cortical responses reflecting melodic expectations [14], a fixed temporal shift of 200 ms was applied to align EEG signals with the corresponding auditory stimuli. Concretely, the starting point of each EEG segment was shifted forward by 200 ms, which corresponds to 25 samples at a sampling rate of 125 Hz. This temporal alignment is consistent with evidence that higher-order auditory cortical responses to music typically emerge with latencies in the range of approximately 150–250 ms [14]. Importantly, this delay value was not tuned in the present study, but was adopted directly from Akama et al. [37], who systematically compared multiple latency conditions and demonstrated that a delay of approximately 200 ms yielded the highest song ID classification accuracy.
@@
EEG normalization was performed independently for each 3-s EEG segment and for each channel. Specifically, a RobustScaler [56] based on the median and interquartile range was applied to the time series of each channel, followed by clamping the scaled values to the range [20, 20][-20,\,20]. This normalization strategy suppresses the influence of transient large-amplitude noise and artifacts while preserving the relative temporal structure of the EEG signals. Because normalization was applied independently to each sample, no statistics were shared across training, validation, or test sets, thereby preventing information leakage. Implementation details are provided in the function normalize_EEG_4 in the file preprocessing_eegmusic_dataset_3s.py.
@@
We constructed three types of artificial neural network (ANN) representations directly from the acoustic signals, as summarized in Figure 9. These representations were designed to capture complementary aspects of auditory processing and were used as teacher signals for masked-prediction-based multitask pretraining. Specifically, we extracted (i) an acoustic representation derived from a self-supervised music foundation model [42], and (ii) predictive representations derived from Surprisal and Entropy computed from an autoregressive music language model [43]. All three ANN representations were discretized into 128 levels (0–127) and served as prediction targets during pretraining.
In the PredANN framework, a central assumption is that ANN representations that closely resemble human cortical representations can serve as effective teacher signals for EEG-based representation learning [37]. Accordingly, we aimed to use ANN representations that are as neurophysiologically plausible as possible. A growing body of prior work has demonstrated that self-supervised learning (SSL) models capture human cortical representations more faithfully than supervised models, particularly in the auditory domain [48, 49, 57, 58]. These findings motivate the use of SSL-based models for extracting teacher representations from audio.
Moreover, prior studies have shown that the architectural properties of ANN models play a critical role in their correspondence with cortical representations. In particular, convolutional layers have been reported to exhibit lower predictive power for brain activity compared with Transformer layers [48]. In contrast, the functional hierarchy formed by Transformer layers has been shown to align well with the hierarchical organization of the human cerebral cortex [48]. Consistent with these observations, multiple studies suggest that Transformer-based models exhibit stronger representational similarity to cortical responses than alternative architectures [48, 49].
Based on this converging evidence, we adopted Transformer-based SSL models to extract ANN representations from audio. Specifically, we used MuQ [42], a masked-prediction-based Transformer model, to obtain acoustic representations, and MusicGen [43], an autoregressive Transformer model, to derive predictive representations in the form of Surprisal and Entropy. The detailed procedures for computing and discretizing these representations are described in the following sections and formalized in dedicated algorithms in Supplementary information.
@@
To construct ANN representations that primarily encode acoustic information, we employed MuQ [42], a Transformer-based masked-prediction self-supervised music foundation model. MuQ has been shown to outperform representative SSL models such as MERT [59] and MusicFM [60] across nearly all tasks in the MARBLE benchmark [61]. Additionally, MuQ-MuLan, which incorporates MuQ as an audio encoder, achieves state-of-the-art performance in zero-shot music tagging, surpassing MuLan [62] and Microsoft-CLAP 2023 [63]. These results indicate that MuQ provides a strong and expressive acoustic representation suitable as a teacher signal.
For each musical piece, continuous MuQ embeddings were extracted from 30-second audio chunks following the original MuQ preprocessing configuration. Audio signals were converted to mono by channel averaging, segmented into 30-second waveforms, and resampled to 24 kHz using a Kaiser-windowed sinc filter, implemented in the librosa package [64] via librosa.resample with res_type="kaiser_fast". MuQ was then applied to each 30-second chunk, and the last-layer hidden representations were extracted, yielding a sequence of frame-wise embeddings at a temporal resolution of 25 Hz (one embedding every 40 ms). Formally, let $\mathcal{T}_{30\mathrm{s}}=\{1,\dots,750\}$ denote the set of time frame indices within a 30-second audio chunk, where the temporal resolution is 25 Hz (40-ms intervals). The resulting continuous acoustic representation $\mathbf{M}_{\mathrm{raw}}$ is defined as:
@@
The use of 24 kHz sampling rate and fixed 30-second input length strictly follows the MuQ pretraining setup [42]. Extracting features under conditions that closely match the pretraining distribution reduces out-of-distribution effects and ensures that the resulting embeddings faithfully reflect the learned acoustic structure.
@@
Thus, the model explicitly distinguishes between continuous acoustic features $\mathbf{m}_{\mathrm{raw}_{t}}$ used as decoder inputs and discrete acoustic tokens $m_{\mathrm{disc}_{t}}$ used as supervision signals, while maintaining a one-to-one temporal correspondence between them. This frame-wise formulation provides the necessary foundation for constructing 3-second MuQ representations aligned with EEG segments in subsequent stages of the model. The detailed procedures for extracting continuous MuQ embeddings and performing K-means discretization are formalized in Algorithm 1 and Algorithm 2.
@@
To construct ANN representations that primarily encode predictive information, we employed MusicGen-large (facebook/musicgen-large) from the Audiocraft framework [43] as an autoregressive model. MusicGen is the first model to achieve high-quality, long-duration, and controllable music generation using a single-stage Transformer language model. It is trained as a language model over discrete audio tokens produced by EnCodec [46], which applies residual vector quantization (RVQ) with multiple codebooks.
@@
A key contribution of MusicGen lies in its treatment of multi-codebook RVQ tokens through codebook interleaving patterns. Copet et al. proposed and evaluated several patterns, including flattening, delay, parallel, and coarse first [43]. While the flattening pattern provides exact autoregressive factorization, the delay pattern uses an inexact but computationally efficient decomposition. Despite reducing computational cost by approximately a factor of four, evaluation showed that the delay pattern achieves similar generation quality to flattening. The MusicGen-large model used in this study is trained with the delay pattern, which we adopt throughout our experiments. We compute predictive features under an empty text condition, corresponding to the model’s unconditional distribution learned through classifier-free guidance training.
@@
The continuous frame-wise Surprisal and Entropy sequences, referred to as Raw Surprisal and Raw Entropy, are used as continuous decoder inputs during masked-prediction pretraining. In parallel, these continuous values are discretized into 128 levels using equal-frequency (quantile) binning to obtain discrete predictive tokens used as ground-truth supervision. Specifically, for each feature type, all values across all songs and segments are pooled to form a one-dimensional set $u$. The global quantile edges are then computed as
$e_{k}=\mathrm{Quantile}(u,k/128),\quad k=0,\dots,128,$
where $e_{k}$ denotes the $k/128$ quantile of $u$. These values define $128$ quantile bins. Each continuous value $u$ is then assigned a bin index $b\in\{0,\dots,127\}$ such that
$e_{b}\leq u<e_{b+1},$
with the maximum value assigned to bin $127$.
By replacing each continuous value with its corresponding discrete token, the resulting discrete sequences (Discretized Surprisal and Discretized Entropy) maintain a strict one-to-one temporal correspondence with the continuous inputs.
The complete procedures for computing frame-wise Surprisal and Entropy and for quantile-based discretization are formalized in Algorithm 3 and Algorithm 4.
@@
Specifically, for each 30-second chunk, MusicGen [43] is applied to compute k1 logits at a temporal resolution of 20 ms, yielding 1500 frames per chunk. Surprisal and Entropy are then computed frame-wise from these logits, and subsequently discretized into 128 bins using the same quantile-based discretization procedure as in the default setting. By construction, this approach prevents MusicGen from exploiting autoregressive context that spans across adjacent 30-second chunks, thereby strictly eliminating potential information leakage from unused acoustic context. The results of this conservative computation are presented in Figure 10.
@@
For MuQ [42], which provides acoustic embeddings at a temporal resolution of 40 ms (25 Hz), the starting frame index is defined as
@@
Recent progress in self-supervised pretraining has established masked modeling with Transformer architectures as a powerful paradigm for representation learning. In natural language processing, BERT [65] demonstrated that predicting masked tokens enables the acquisition of rich bidirectional contextual representations, while the GPT series showed that autoregressive pretraining effectively captures long-range dependencies [66, 67]. In computer vision, this idea was reformulated as masked image modeling, where approaches such as MAE [68], BEiT [69], and SimMIM [70] revealed that reconstructing missing content from partial observations leads to highly transferable representations.
A key step in transferring these ideas to EEG representation learning is exemplified by LaBraM [71]. Inspired by masked modeling in vision, it reformulates EEG signals as collections of time–channel patches and applies Transformer-based masked reconstruction to large-scale EEG data. Through architectural choices such as temporal patch encoders, learnable channel and time embeddings, and Pre-Norm self-attention, LaBraM demonstrated that masked modeling is an effective and scalable principle for EEG representation learning.
The original PredANN framework, by contrast, was built around a CLIP [72]-style contrastive learning objective using CNN-based encoders optimized with an InfoNCE loss. While this design successfully aligned EEG representations with stimulus-derived ANN representations, it fundamentally differs from the generative and structural learning paradigm underlying recent masked autoencoding approaches. Motivated by the success of LaBraM and masked modeling in other modalities, we therefore redesign PredANN into a Transformer-based masked modeling framework.
Specifically, we propose PredANN++, which extends PredANN by replacing the CLIP-style CNN encoder with a masked autoencoding Transformer architecture. Rather than discretizing EEG signals themselves as in LaBraM, PredANN++ treats discretized stimulus-derived representations (MuQ [42], Surprisal, and Entropy) as teacher sequences and formulates pretraining as masked prediction of these discrete targets. This design preserves the original PredANN philosophy of learning EEG representations aligned with stimulus-level ANN representations, while adopting the masked generative learning paradigm shown to be effective in LaBraM. The overall architecture is illustrated in Fig. 8 and is formalized below.